# Emojis Decoded: Leveraging ChatGPT for Enhanced Understanding in Social Media Communications

### v2

@@ Abstract
Emojis, which encapsulate semantics beyond mere words or phrases, have become prevalent in social network communications. This has spurred increasing scholarly interest in exploring their attributes and functionalities. However, emoji-related research and application face two primary challenges. First, researchers typically rely on crowd-sourcing to annotate emojis in order to understand their sentiments, usage intentions, and semantic meanings. Second, subjective interpretations by users can often lead to misunderstandings of emojis and cause the communication barrier. Large Language Models (LLMs) have achieved significant success in various annotation tasks, with ChatGPT demonstrating expertise across multiple domains. In our study, we assess ChatGPT’s effectiveness in handling previously annotated and downstream tasks. Our objective is to validate the hypothesis that ChatGPT can serve as a viable alternative to human annotators in emoji research and that its ability to explain emoji meanings can enhance clarity and transparency in online communications. Our findings indicate that ChatGPT has extensive knowledge of emojis. It is adept at elucidating the meaning of emojis across various application scenarios and demonstrates the potential to replace human annotators in a range of tasks.
@@ 1 Introduction
Emojis, as prevalent non-verbal tokens in social networks, are crucial in developing an effective discussion in online communication. These language units encode rich semantic meanings that range from emotions, objects, to actions. The crucial role of emojis has attracted a number of researchers to study various emoji-related topics, from emoji functionality, emoji usage patterns, to the emoji application on downstream tasks. However, there are two challenges for emoji research and emoji applications for social users. For emoji research, to determine the meaning, sentiments, or intentions of emojis in different scenarios, researchers usually hire human annotators to carry out annotation jobs (Lu et al. 2018; Hu et al. 2017), which is not scalable and time consuming. For social users, their differences in demographics, cultural background, or personal experience could lead to misunderstanding of emoji, such as [Uncaptioned image] (upside-down face) to express irony, and incorrect interpretation may cause miscommunication and block emoji diffusion.
Large language models (LLMs), pre-training on large corpuses, have demonstrated the remarkable achievements in multiple natural language processing tasks. Among various LLMs, ChatGPT (GPT3.5 and GPT4) (Ouyang et al. 2022; OpenAI 2023), has been utilized on downstream application tasks in different domains due to its inexpensive costs and human-aligned response (Ouyang et al. 2022). There is an increasing interest in the use of ChatGPT on social science research questions, especially questions about social networks (Ziems et al. 2023; Zhu et al. 2023). ChatGPT could be the potential solution to these two challenges. First, motivated by the superiority of ChatGPT over human annotators in various text annotation tasks (Kocoń et al. 2023), we hypothesize that ChatGPT has a deep understanding of emoji applications and can replace human workers to label emoji attributes. Second, if ChatGPT has a similar understanding of emojis to humans, when users encounter unfamiliar emojis, they can ask ChatGPT to explain the meaning of the emoji to increase information transparency. On the other hand, previous work suggests that there exist hallucinations in LLMs’ generations (Zhang et al. 2023), inconsistencies between the model’s output and the real-world facts or user inputs. Whether ChatGPT can produce generations similar to humans for the meaning of emojis is still questionable.
To verify ChatGPT’s potential on emoji research and application, our paper conducts a qualitative and quantitative study about ChatGPT’s understanding on emoji usage in multiple tasks. Based on the current work on emoji studies, we propose three research questions to probe the ChatGPT’s capability on emoji-related tasks.
@@
To answer these research questions, we first ask ChatGPT to explain the semantics, sentiments, and intentions of emojis with or without text context, across different platforms or cultures (Section 3, 4, 5). We find that the interpretation of ChatGPT on emojis is consistent with the human answers in the existing literature in most cases. Next, we ask ChatGPT to generate the associated emoji patterns with multiple communities (gender, platform, hashtag, and culture) to probe ChatGPT’s knowledge about emoji usage in different communities (Section 6). Finally, we test ChatGPT on emoji-related downstream tasks, i.e. irony annotation and emoji prediction (Section 7, 8), and the results reveal that ChatGPT is capable of using the emoji information for irony annotation and making emoji recommendations based on user identity.
@@ 2.2 LLMs in Computational Social Science
With the prevalence of LLMs, there are emerging researches evaluating their performances on various tasks (Chang et al. 2023), including Computational Social Science (CSS) tasks (Ziems et al. 2023; Zhu et al. 2023). Researchers have examined ChatGPT’s capacities in the zero-shot or few-shot (with in-context learning) settings, in various text annotation tasks such as political affiliation classification of tweets (Törnberg 2023); hate speech detection (Huang, Kwak, and An 2023; Li et al. 2023b; Zhu et al. 2023); discourse acts (Ostyakova et al. 2023); sentiment analysis and bot detection(Zhu et al. 2023; Zhou et al. 2023b). A comprehensive evaluation (Ziems et al. 2023) shows ChatGPT exhibits unsatisfactory results on tasks that have complex structure or have subjective taxonomies whose semantics differ from definitions learned in pretraining, while achieving better results on tasks that have objective ground truth (e.g., misinformation and fact checking) or have labels with explicit colloquial definitions in the pretraining data (e.g., emotion and political stance). Nevertheless, the results demonstrate the potential of using LLMs to help researchers with CSS tasks and general users with social-related daily activities.
Not too many studies focus on emojis and ChatGPT at the same time, and emoji as an emerging “language” has not received enough attention in LLM studies. Only a few works in the literature suggest the ability of ChatGPT in emoji understanding, but do not provide detailed evidence and discussion. de Janeiro (2023) asks ChatGPT to generate different facial emojis to represent the valence of the words and Peng et al. (2023) translates natural language to emoji sequence. The most similar works on the use of ChatGPT in emoji-related tasks are Belal, She, and Wong (2023) and Kocoń et al. (2023). Belal, She, and Wong (2023) finds that ChatGPT performs better than lexicon-based approaches for the sentiment prediction task on a soccer tweet dataset with emojis, and they imply the emoji understanding ability of ChatGPT. In Kocoń et al. (2023), the emoji prediction task, as one of the evaluation tasks in their work, reveals the ChatGPT’s performance on emoji recommendation.
In comparison, we thoroughly examine the consistency of ChatGPT’s annotations on emoji semantics, sentiment, and intentions with humans and probe ChatGPT’s knowledge about the emoji usage patterns linked to various communities. We evaluate ChatGPT’s performance on emoji-related tasks and explore the importance of emojis to ChatGPT in downstream tasks. Our evaluation task is performed across multiple languages and multiple platforms, which qualitatively and quantitatively demonstrate the ChatGPT’s ability on emoji understandings in broader contexts.
@@ 3.1 Emoji Semantics without Context
where [Emoji] is the Unicode of selected emojis. To compare the understandings between ChatGPT and humans, we also collect the meaning of emojis on the Emojipedia website, [^2] https://emojipedia.org/meanings which is researched and written by Emojipedia editors and lexicographers, as human annotations of emoji meanings. We present three representative examples for the emoji [Uncaptioned image] (person shrugging), [Uncaptioned image] (purple heart), [Uncaptioned image] (blue heart) in Figure 1, highlighting the same contents mentioned by humans and ChatGPT by green and different contents by red color.
@@ 3.2 Emoji Semantics with Text Context
The accurate explanation of ChatGPT in emoji semantics without text context follows our expectation due to the notable performance of ChatGPT in sentiment analysis and tweet related tasks (Zhu et al. 2023; Törnberg 2023). But it is well known that emojis will evolve extensive semantics according to the applied text context. For example, [Uncaptioned image] (goat) can mean the greatest person, and [Uncaptioned image] (snake) can mean Python language. We thus explore whether ChatGPT can interpret an emoji’s meaning when giving tweets containing emojis with extensive semantics. We choose a few emojis ([Uncaptioned image] goat, [Uncaptioned image] honey bee, and [Uncaptioned image] snake) with widely accepted special meanings 333selected from this website https://bestlifeonline.com/emoji-meanings/ and randomly extract tweets in 2022 that contain these emojis via Twitter API. We feed these tweets to ChatGPT and ask ChatGPT to give the semantic explanation based on the context of the tweet. We present the prompt and the GPT4 response in Figure 2.
Figure 2 illustrates that ChatGPT’s interpretations of emoji semantics align closely with human understanding in most instances. ChatGPT accurately interprets the emoji [Uncaptioned image] (goat) as “the greatest of all time.” Additionally, it interprets [Uncaptioned image] in a Valentine’s context as “be mine.” In the third example featuring the [Uncaptioned image] emoji, ChatGPT recognizes its use in symbolizing the Python programming language. However, in the last example with the tweet context “Whole lotta snakes in my grass [Uncaptioned image][Uncaptioned image],” we argue that the emoji should symbolize the animal snake rather than untrustworthy people. The frequent use of the snake emoji to represent betrayal might have led to this misinterpretation. In conclusion, ChatGPT generally demonstrates a precise understanding of emoji semantics in textual contexts. However, when an emoji represents a less commonly used object, ChatGPT may provide erroneous interpretations.
@@ 3.3 Emoji Semantics with Language/Cultural Context
In addition to the application in the English context, emojis also show prevalence in other specific language and culture contexts (Chen et al. 2019), and explaining emoji semantics in the Chinese context is more challenging. With the similar experiment setup in Section 3.2, we first choose emojis that have unique meanings on Chinese social networks and test if ChatGPT can properly explain their semantics, that is, [Uncaptioned image] (slightly smiling face), [Uncaptioned image] (pill), and [Uncaptioned image] (doge). Additionally, we include an emoji combo, [Uncaptioned image] [Uncaptioned image], that consists of a cow face and a beer mug. We prompt ChatGPT to explain the usage of these emojis. The prompts are written in English or Chinese, with or without specification of usage in Chinese pop culture.
We find that GPT models can generate desired interpretations in all cases, but only with appropriate prompts. The prompts need to include very specific information sometimes. For instance, when prompted with “Explain [Emoji]”, GPT models failed to generate desired explanations and answered with literal interpretations from their Unicode names or visual cues (e.g., [Uncaptioned image] as friendly instead of its common ironical use in Chinese pop culture (Wang 2022)). When prompted in Chinese, only [Uncaptioned image] [Uncaptioned image] got the correct explanation. However, [Uncaptioned image] and [Uncaptioned image] get desired explanations when prompted with “Explain the [Emoji] in the context of Chinese pop culture”, no matter whether prompted in English or Chinese. For [Uncaptioned image], we had to specifically mention its homophonic usage with yào wán through a multi-turn interaction with ChatGPT.
@@
ChatGPT can output consistent explanations on emoji semantics when giving no context or specific English context. However, when probing the special meaning of an emoji in Chinese context, ChatGPT are prone to hallucinations with English prompts and more accurate explanation when prompting in Chinese.
@@ 4 Emoji Sentiment
Expressing sentiment is one of the main functionalities of using emojis (Ai et al. 2017), and for most emojis, especially face emojis. Previous work has performed human annotation of emoji sentiment (Hu et al. 2017) and incorporated emoji information to improve the downstream sentiment (Felbo et al. 2017; Chen et al. 2019). In this section, we ask ChatGPT to annotate the emoji sentiment and evaluate ChatGPT performance on the sentiment of tweets with emojis.
@@ 4.1 Emoji Sentiment Annotation
We first ask ChatGPT to label the 50 most frequently used emojis in Section 3.1 with positive, neutral, and negative sentiment and compare the differences between human and ChatGPT annotations. We present the prompt details in Table 8(a) in Appendix.
We compare the GPT4 annotation with the human annotation in Hu et al. (2017), and the only difference is the sentiment annotation of emoji [Uncaptioned image] flushed face. The human annotation is neutral and the response of ChatGPT is negative, considering feelings of embarrassment or a sense of being overwhelmed in [Uncaptioned image]. The same sentiment annotation indicates that ChatGPT can fully understand the sentiment in the emojis. From the response from ChatGPT, it often mentions that it requires the text context to determine the emoji sentiment, so we discuss the influence of emojis for ChatGPT on the downstream sentiment analysis task.
@@ 4.2 Sentiment Prediction with Emojis
Previous researchers have developed multiple methods to involve emoji information to enhance sentiment classification (Chen et al. 2019; Felbo et al. 2017). We also ask: When ChatGPT performs the sentiment classification task, does it consider the emoji in the texts in its predictions? For the experiment setup, we choose the sentiment analysis dataset from the TweetEval benchmark (Barbieri et al. 2020) with three labels: negative, neutral, and positive. We extract tweets with emojis from the data and obtain 805 tweets. We perform sentiment prediction by in-context learning (ICL) with GPT4 and GPT3.5 with 5 demonstrations on these tweets and obtain 70.31% and 71.67% accuracy for GPT4 and GPT3.5, respectively. The fine-tuning Roberta model with pretraining on the tweet corpus can achieve 69.3% accuracy (Barbieri et al. 2020), which indicates a reliable sentiment prediction from ChatGPT.
Next, we remove the emojis from the input text and ask ChatGPT to assign a sentiment label again, and the performance accuracy decreases to 69.56% and 70.19% for GPT4 and GPT3.5, respectively. However, the accuracy decrease cannot demonstrate that ChatGPT utilizes emoji information for emoji prediction, since the accuracy change is not significant and the emoji removal may also change the ground truth sentiment for the text. We compare the sentences in which ChatGPT changes the sentiment prediction after removing the emojis and find that 19.50% / 18.13% tweets flip the prediction labels for GPT3.5 / GPT4 annotations. We present three randomly selected tweets with flipped predictions for GPT4 in Table 1.
@@
GPT4v and GPT4 annotate the similar emoji sentiment labels with humans, and when performing the sentiment classification task, ChatGPT considers emoji sentiment to enhance the prediction.

### v3

@@ Abstract
Emojis, which encapsulate semantics beyond words or phrases, have become prevalent in social network communications. This has spurred increasing scholarly interest in exploring their attributes and functionalities. However, emoji-related research and application face two primary challenges. First, researchers typically rely on crowd-sourcing to annotate emojis in order to understand their sentiments, usage intentions, and semantic meanings. Second, subjective interpretations by users can often lead to misunderstandings of emojis and cause a communication barrier. Large Language Models (LLMs) have achieved significant success in various annotation tasks, with ChatGPT demonstrating expertise across multiple domains. In our study, we assess ChatGPT’s effectiveness in handling previously emoji-annotated and downstream tasks. Our objective is to validate the hypothesis that ChatGPT can serve as an alternative to human annotators in emoji research and that its ability to explain emoji meanings can enhance clarity and transparency in online communications. Our findings indicate that ChatGPT has extensive knowledge of emojis. It is adept at explaining the meaning of emojis across various application scenarios and demonstrates the potential to replace human annotators in a range of tasks.
@@ 1 Introduction
Emojis, as prevalent non-verbal tokens in social networks, are crucial in developing an effective discussion in online communication. These language units encode rich semantic meanings that range from emotions, objects, to actions. The crucial role of emojis has attracted a number of researchers to study various emoji-related topics, from emoji functionality, emoji usage patterns, to the emoji application on downstream tasks. However, there are two challenges for emoji research and emoji applications for social users. For emoji research, to determine the meaning, sentiments, or intentions of emojis in different scenarios, researchers usually hire human annotators to carry out annotation jobs (Lu et al. 2018; Hu et al. 2017), which is not scalable and time consuming. For social users, their differences in demographics, cultural background, or personal experience could lead to misunderstanding of emoji, such as [Uncaptioned image] (upside-down face) to express irony, and incorrect interpretation may cause miscommunication and block emoji diffusion (Leonardi 2022).
We hypothesize that ChatGPT could offer potential solutions to two key challenges related to emoji interpretation, based on prior explorations. First, given ChatGPT’s demonstrated superiority over human annotators in various text annotation tasks (Kocoń et al. 2023), we believe it may possess a deep understanding of emoji usage. As such, it could replace human workers in labeling emoji attributes, reducing the need for crowd-sourcing efforts and lowering costs. Second, when users encounter unfamiliar emojis or are unsure of their meanings, ChatGPT’s ability to explain emoji semantics could effectively prevent misunderstandings.
On the other hand, previous work suggests that there exist hallucinations in LLMs’ generations (Zhang et al. 2023), inconsistencies between the model’s output and the real-world facts or user input. Whether ChatGPT can produce generations similar to humans for the meaning of emojis is still questionable. Delineating the boundaries of ChatGPT’s capabilities across emoji-related tasks of varying difficulty is essential for driving progress in future emoji research.
To verify ChatGPT’s potential on emoji research and application, our paper conducts a set of qualitative and quantitative studies about ChatGPT’s understanding on emoji usage in multiple tasks. Based on the current work on emoji studies, we propose three research questions to probe the ChatGPT’s capability on emoji-related tasks.
@@
To answer these research questions, we first ask ChatGPT to explain the semantics, sentiments, and intentions of emojis with or without text context, across different platforms or cultures (Section 3, 4, 5). We find that the interpretation of ChatGPT on emojis is consistent with human responses in the existing literature in most cases. Next, we ask ChatGPT to generate the associated emoji patterns with multiple communities (gender, platform, hashtag, and culture) to probe ChatGPT’s knowledge about emoji usage in different communities (Section 6). Finally, we test ChatGPT on emoji-related downstream tasks, i.e., irony annotation and emoji prediction (Section 7, 8), and the results reveal that ChatGPT is capable of using the emoji information for irony annotation and making emoji recommendations based on user identity.
@@ 2.2 LLMs in Computational Social Science
With the prevalence of LLMs, there are emerging researches evaluating their performances on various tasks (Chang et al. 2023; Liu et al. 2024), including Computational Social Science (CSS) tasks (Ziems et al. 2023; Zhu et al. 2023). Researchers have examined ChatGPT’s capacities in the zero-shot or few-shot (with in-context learning) settings, in various text annotation tasks such as political affiliation classification of tweets (Törnberg 2023); hate speech detection (Huang, Kwak, and An 2023; Li et al. 2023b; Zhu et al. 2023); discourse acts (Ostyakova et al. 2023); instruction quality evaluation (Xu et al. 2024); sentiment analysis and bot detection(Zhu et al. 2023; Zhou et al. 2023b). A comprehensive evaluation (Ziems et al. 2023) shows ChatGPT exhibits unsatisfactory results on tasks that have complex structure or have subjective taxonomies whose semantics differ from definitions learned in pretraining, while achieving better results on tasks that have objective ground truth (e.g., misinformation and fact checking) or have labels with explicit colloquial definitions in the pretraining data (e.g., emotion and political stance). Nevertheless, the results demonstrate the potential of using LLMs to help researchers with CSS tasks and general users with social-related daily activities.
Despite the growing prominence of emojis as an emerging “language,” limited studies focus on the intersection of emojis and ChatGPT. Only a few works suggest ChatGPT’s ability to comprehend emojis. de Janeiro (2023) examines ChatGPT’s ability to generate various facial emojis to represent the valence of words, while Peng et al. (2023) explores translating natural language into emoji sequences. Belal, She, and Wong (2023) and Kocoń et al. (2023) find that ChatGPT can interpret emojis in texts to enhance sentiment prediction. The most relevant concurrent work is by Lyu et al. (2024), which investigates whether GPT-4V understands emoji interpretations and correctly employs emojis when given social media posts. However, their study does not cover the functionality or intent of emojis nor discuss emoji usage across different communities.
In contrast, we thoroughly examine the consistency of ChatGPT’s annotations on emoji semantics, sentiment, and intentions with human interpretations and probe ChatGPT’s knowledge of emoji usage patterns across various communities. We evaluate ChatGPT’s performance on emoji-related downstream tasks and explore the significance of emojis to ChatGPT in these tasks. Our evaluation spans multiple languages and platforms, qualitatively and quantitatively demonstrating ChatGPT’s understanding of emojis in broader contexts.
@@ 3.1 Emoji Semantics without Context
where [Emoji] is the Unicode of selected emojis. To compare the understandings between ChatGPT and humans, we also collect the meaning of emojis on the Emojipedia website, [^2] https://emojipedia.org/meanings which is researched and written by Emojipedia editors and lexicographers, as human annotations of emoji meanings. Two authors of the paper independently compare the human annotations and ChatGPT responses, finding that ChatGPT and humans provide similar explanations for all 50 emojis. However, we note that both ChatGPT and human annotators highlight points that the other may not cover.
We present three representative examples for the emoji [Uncaptioned image] (person shrugging), [Uncaptioned image] (purple heart), [Uncaptioned image] (blue heart) in Figure 1, highlighting the same contents mentioned by humans and ChatGPT by green and different contents by red color.
@@ 3.2 Emoji Semantics with Text Context
The accurate explanation of ChatGPT in emoji semantics without text context follows our expectation due to the notable performance of ChatGPT in sentiment analysis and tweet-related tasks (Zhu et al. 2023; Törnberg 2023). However, it is well known that emojis will evolve extensive semantics according to the applied text context. For example, [Uncaptioned image] (goat) can mean the greatest person, and [Uncaptioned image] (snake) can mean the Python language in the specific context. We thus explore whether ChatGPT can interpret an emoji’s meaning when giving tweets containing emojis with extensive semantics. We choose a few emojis ([Uncaptioned image] goat, [Uncaptioned image] honey bee, and [Uncaptioned image] snake) with widely accepted special meanings333selected from this website https://bestlifeonline.com/emoji-meanings/ and randomly extract tweets in 2022 that contain these emojis via Twitter API. We feed these tweets to ChatGPT and ask ChatGPT to give the semantic explanation based on the context of the tweet.
We present the prompt and the GPT4 response in Figure 2, which illustrates that ChatGPT’s interpretations of emoji semantics align closely with human understanding in most instances. ChatGPT accurately interprets the emoji [Uncaptioned image] (goat) as “the greatest of all time.” Additionally, it interprets [Uncaptioned image] in a Valentine’s context as “be mine.” In the third example featuring the [Uncaptioned image] emoji, ChatGPT recognizes its use in symbolizing the Python programming language. However, in the last example with the tweet context “Whole lotta snakes in my grass [Uncaptioned image][Uncaptioned image],” we argue that the emoji should symbolize the animal snake rather than untrustworthy people. The frequent use of the snake emoji to represent betrayal might have led to this misinterpretation. In conclusion, ChatGPT generally demonstrates a precise understanding of emoji semantics in textual contexts. However, when an emoji represents a less commonly used object, ChatGPT may provide erroneous interpretations.
@@ 3.3 Emoji Semantics with Language/Cultural Context
In addition to the application in the English context, emojis also show prevalence in other specific language and culture contexts and emojis may evolve to some unique meanings in a specific language context (Chen et al. 2019; Lu et al. 2016). ChatGPT, as a multilingual language model, has demonstrated strong cross-linguistic understanding abilities (OpenAI 2023). To evaluate the feasibility of ChatGPT replacing human annotators in different language contexts, we select a few emojis with unique meanings in the Chinese context and observe whether ChatGPT could correctly capture the specific semantics.
With a similar experiment setup as in Section 3.2, we first choose emojis that have unique meanings on Chinese social networks and test if ChatGPT can properly explain their semantics, that is, [Uncaptioned image] (slightly smiling face), [Uncaptioned image] (doge), [Uncaptioned image] (pill), [Uncaptioned image] (bug), [Uncaptioned image](lemon), [Uncaptioned image] (melon), and [Uncaptioned image] (bird). In the Chinese pop culture context, [Uncaptioned image] ironically expresses dislike and ridicule; [Uncaptioned image] conveys a sense of irony or sarcasm; [Uncaptioned image] and [Uncaptioned image] use a homophone pun, meaning “it’s over” and expressing support, respectively; [Uncaptioned image] express envy and jealousy; [Uncaptioned image] means gossip or drama; [Uncaptioned image] refers to not showing up for a planned meeting, when it’s used as a pigeon.
Additionally, we include an emoji combo, [Uncaptioned image] [Uncaptioned image], that consists of a cow face and a beer mug (a homophone pun to express “so cool”). We prompt ChatGPT to explain the usage of these emojis. The prompts are written in English or Chinese, with or without specification of usage in Chinese pop culture. For some emojis, we include appropriate hints, such as connections to homophone puns.
We find that ChatGPT can generate desired interpretations in all cases, but only with appropriate prompts. The prompts need to include very specific information sometimes. For instance, when prompted with “Explain [Emoji]”, GPT models failed to generate desired explanations and answered with literal interpretations from their Unicode names or visual cues (e.g., [Uncaptioned image] as friendly instead of its common ironical use in Chinese pop culture (Wang 2022)). When prompted in Chinese, only [Uncaptioned image] [Uncaptioned image] got the correct explanation. However, [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image] get desired explanations when prompted with “Explain the [Emoji] in the context of Chinese pop culture”, no matter whether prompted in English or Chinese. For [Uncaptioned image] and [Uncaptioned image], we had to specifically mention its homophonic usage with yào wán and chōng, respectively. Similarly for [Uncaptioned image] we had to mention its usage as a pigeon.
@@
In summary, ChatGPT is capable of correctly explaining the emojis in the Chinese pop culture context. However, prompting in Chinese or mentioning appropriate hints is necessary to better elicit such knowledge in some cases.
@@
ChatGPT can output consistent explanations on emoji semantics when giving no context or specific English context. However, when probing the special meaning of an emoji in the Chinese pop culture context, ChatGPT is prone to hallucinations with English prompts but more accurate explanations when prompted in Chinese or given appropriate hints.
@@ 4 Emoji Sentiment
Expressing sentiment is one of the primary functions of emojis, especially face emojis (Ai et al. 2017). Previous research has employed human annotation to investigate emoji sentiment (Hu et al. 2017) and has incorporated emoji information to enhance downstream sentiment analysis (Felbo et al. 2017; Chen et al. 2019). In this section, we prompt ChatGPT to annotate emoji sentiment and compare its annotations with those of humans. Additionally, we evaluate ChatGPT’s performance in predicting the sentiment of tweets containing emojis for downstream tasks.
@@ 4.1 Emoji Sentiment Annotation
We first ask GPT4 to label the 50 most frequently used emojis in Section 3.1 with positive, neutral and negative sentiment with the prompt in Table 12(a) in Appendix.
Compared to the human annotations in Hu et al. (2017), we observe that ChatGPT and human annotators share the same sentiment annotation for almost all emojis (49 out of 50). The only difference is the annotation for the [Uncaptioned image] (flushed face) emoji. While the human annotation is neutral, ChatGPT’s response is negative, considering feelings of embarrassment or a sense of being overwhelmed associated with [Uncaptioned image]. This discrepancy indicates that ChatGPT can interpret the sentiment of emojis but may differ from human annotations in specific cases. ChatGPT’s response often notes that it requires text context to accurately determine emoji sentiment. Therefore, we discuss the influence of emojis on ChatGPT’s performance in downstream sentiment analysis tasks.
@@ 4.2 Sentiment Prediction with Emojis
Previous researchers have emphasized the importance of emoji information to judge sentiment in the given inputs (Chen et al. 2019; Felbo et al. 2017). We also investigate whether ChatGPT considers emojis in its sentiment classification predictions. For the experiment setup, we choose the sentiment analysis dataset from the TweetEval benchmark (Barbieri et al. 2020), which includes three labels: negative, neutral, and positive. We extract tweets containing emojis from the dataset, resulting in 805 tweets. We perform sentiment prediction using in-context learning (ICL) with GPT4 and GPT3.5, providing 5 demonstrations for each model. The models achieve 70.31% and 71.67% accuracy for GPT4 and GPT3.5, respectively. The fine-tuned RoBERTa model, pretrained on the tweet corpus, achieves 69.3% accuracy (Barbieri et al. 2020), indicating that ChatGPT provides reliable sentiment predictions.
Next, we remove the emojis from the input text and ask ChatGPT to assign a sentiment label again. The performance accuracy decreases to 69.56% and 70.19% for GPT4 and GPT3.5, respectively. However, this decrease in accuracy does not conclusively demonstrate that ChatGPT utilizes emoji information for sentiment prediction, as the change is not significant and removing emojis may also alter the ground-truth sentiment of the text. We compare the sentences where ChatGPT changes the sentiment prediction after removing the emojis and find that 19.50% and 18.13% of tweets flip the prediction labels for GPT3.5 and GPT4 annotations, respectively. We present three randomly selected tweets with flipped predictions for GPT4 in Table 1.
@@
GPT-4 annotates emoji sentiment labels similarly to humans and considers emoji sentiment to enhance its predictions when performing sentiment classification tasks.